OpenStudy (anonymous):

simplify please :) -(2a - 3b + 1)

OpenStudy (anonymous):

just change the negatives to positives and vice versa. -2a + 3b - 1

OpenStudy (anonymous):

-(2a-3b+1) Multiply -1 by each term inside the parentheses. -2a+3b-

OpenStudy (vishweshshrimali5):

well its already simplified \[-(2a - 3b + 1) = -2a + 3b -1 = 3b - 2a - 1 = 3b - (2a+1)\]

OpenStudy (anonymous):

true

OpenStudy (anonymous):

dam easier then I thought I must be tired, Thanks !